OpenStudy (anonymous):

Write the quadratic function with the following transformations: vertical shrink by 3, left 4, down 9

OpenStudy (mathmale):

The appropriate equation for the quadratic function before any transformation, shrinkage or stretching is simply y=x^2. If the vertex of this parabolic graph is moved h units to the right, then the equation becomes y=(x-h)^2. If the vertex of this latest parabolic graph is now also shifted up by k units, the equation becomes y=(x-h)^2 + k We'll take care of the vertical shrinking later. For now, please use the background info I've given you and the numerical values given in the problem statement to write the equation of this parabola.